OpenStudy (anonymous):

help me please. 3x+4=2x-5 3x+4-4=2x-5-4 3x+2x=-6 5x=-6 5x/5=-6/5=?

OpenStudy (anonymous):

Can anyone help?

OpenStudy (anonymous):

what is the question

OpenStudy (anonymous):

What do I do now?

OpenStudy (anonymous):

R u there

OpenStudy (anonymous):

You have a mistake in the third line already. I'll help.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

After your second line, you should end up with: 3x=2x-9 Now, you need to subtract 2x from both sides of the equation to get rid of the 2x on the right side. 3x-2x=2x-9-2x

OpenStudy (anonymous):

Which leaves you with x=-9

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

The shortcut is to move the term to the other side of the equation, but you have to change the sign. In 3x=2x-9, you can move the 2x term to the left side of the equation, but you need to change it from a + to a -. 3x-2x=-9 x=-9

OpenStudy (anonymous):

You're welcome. Take care.

OpenStudy (anonymous):

ok you 2

OpenStudy (anonymous):

Can you help me with 6(x-4)=3(3x+5) 6

OpenStudy (anonymous):

6x-24=9x+15 -24+15=9x-6x -9=3x -3=x Is this correct?

OpenStudy (anonymous):

did i do this right

OpenStudy (anonymous):

Nope. Remember when you move a term from one side of an equation to the other, you need to change the sign.

OpenStudy (anonymous):

So, you're second line should be: -24-15=9x-6x <---- a negative before the 15

OpenStudy (anonymous):

6x-24 = 9x+15 -24-15 = 9x-6x -39 = 3x x = -39/3 x = -13